OpenStudy (anonymous):

John shot 67 baskets that were worth 2 or 3 points each. He made a total of 160 points. How many of each type of basket did John make? Hint* Set up a system of equations.

OpenStudy (anonymous):

x = 2 points y = 3 points x + y = 67 2x + 3y = 160 x = 67 - y 134 -2y + 3y = 160 y = 26 x = 41